How many unit cells are present in a cube shaped ideal crystal of NaCl of mass 1g?

Grade:12th Pass

3 Answers

Beenu Mathew
7 years ago
HI,
we all know the equation 
d= Z * M/a3 * Na
Now,
the twist....are you ready
d* a3 = Z * M/Na
what is d * a3 ….well....IT IS THE MASS OF 1 UNIT CELL
SO,
now...
we know NaCl has a mass of 35.5g
and Z=4 
so,
mass of 1 unit cell= 4*35.5/Na
                           = 142/Na
Number of unit cells = 10g* Na/142
                              finally you will get an answer of approx 
                              4.2 * 1022 unit cells
CHEERS!!
GREAT QUESTION
VV IMP
BEST OF LUCK!!
 
 
Ranjan Choudhary
6 years ago
using eqn,
d= Z * M/a3 * Na 
or 
d* a3 = Z * M/Na 
 d * a3 …..IT IS THE MASS OF 1 UNIT CELL
One unit cell of NaCl contains 4 NaCl units (ie) 
we know NaCl has a mass of 58.5g
 
mass of 1 unit cell= 4*58.5/Na
                           = 234/Na
Number of unit cells = 1g* Na/234=2.57 *10^212.57×10212.57×102
 
Arbaz khan
5 years ago
Molar mass of NaCl=58.5
We knw in a unit cell we have 4 formula mass of  NaCl, so z=4
d=z*M/a^3*Na
d*a^3=z*M/Na
We know mass=density *vol. => m=d*a^3
This mass represnts mass of 1 unit cell. 
Then unit cell mass=4*58.5/N
Then unit cell in 1g of NaCl crystal=mass of
 
crystal/mass of unit cell
=1/4*58.5*Na => 6*10^23/234 = 2.57*1021 ans.